# Formation of GW230529 from Isolated Binary Evolution and Its Electromagnetic Counterparts

### v1

@@ 3.1 Tidal Disruption Probability
Figure 3 shows the parameter space where the NS can be tidally disrupted using Equation (1). For a given $\chi_{{\rm BH},z}$, BHNS mergers composed of a low-mass BH component and a low-mass NS component are more easily to have tidal disruption. The mass space that allows NS tidal disruption expands significantly with increasing $\chi_{{\rm BH},z}$ and the adoption of a stiffer EoS. We display the 90% credible posterior distributions and the medians of component masses, as well as the medians of BH aligned-spins $\chi_{{\rm BH},z}$, for GW200105, GW200115, and GW230529 (Abbott et al., 2021a; The LIGO Scientific Collaboration et al., 2024) in Figure 3. The explicit results of the tidal disruption probabilities for these three BHNS mergers are listed in Table 2. It is obvious that GW200105 is unlikely to make tidal disruption to generate any EM counterparts, because its BH and NS components are massive with the mass distribution significantly outside the tidal disruption region. Although GW200115 has BH and NS masses much lighter than those of GW200105, tidal disruption can occur with only a very low probability of $P_{\rm tidal}=2.76\%$ under the EoS of DD2. Compared with GW200105 and GW200115 whose BH masses are $\gtrsim 5\,M_{\odot}$, GW230529, as a mgBHNS merger, occupies a larger portion of its mass parameter space located within the disruption region, especially when considering a stiffer EoS. More specifically, the tidal disruption probabilities are $12.8\%$ and $63.2\%$ by adopting the EoS of AP4 and DD2, respectively. The former value is similar to that reported in The LIGO Scientific Collaboration et al. (2024), which was obtained by marginalizing over the EoS.
@@ 3.4 Implication of EM detectability for BHNS Mergers
Total BHNS $32.8\%$ $3.34$ $0.05$ $0.69$ $3.28$ $0.61$ $0.94$
DD2 Mass-gap BHNS $45.0\%$ $4.58$ $0.27$ $3.59$ $4.55$ $0.91$ $1.41$
High-mass BHNS $3.0\%$ $0.31$ 00 $0.13$ $0.29$ $0.02$ $0.03$
Total BHNS $48.0\%$ $4.89$ $0.27$ $3.72$ $4.84$ $0.93$ $1.44$
@@
Based on our population synthesis simulation results, we then simulate the tidal disruption probability, kilonova detectable rate, and GRB detection rate for GW BHNS mergers detected in the near future GW observing era (e.g., O4b and O5). We conservatively consider BHNS mergers occurring within $300\,{\rm Mpc}$, which is the observed distance of GW200105 and GW200115, as well as the representative distance of detectable BHNS mergers in O4 (Abbott et al., 2018b; Zhu et al., 2021b; Colombo et al., 2023; Gupta et al., 2023). We find that $100\%$ and $\gtrsim 90\%$ disrupted BHNS mergers are expected to originate from mgBHNS mergers by adopting the EoS of AP4 and DD2, respectively. When the detection depth exceeds $24\,{\rm mag}$, the survey projects can cover the majority of kilonovae associated with BHNS mergers. However, considering the rapid rise and decline of kilonovae, their signals may be challenging to be identified. Additionally, we find that the detection rate of GRBs originating from mgBHNS mergers is $\sim 1\,{\rm yr}^{-1}$. Therefore, in the future, it may be possible to search for kilonova signals via the target-of-opportunity follow-up observations of both GW and GRB triggers, like the multimessenger observations between GW170817/GRB170817A/AT2017gfo.
@@ 4 Conclusion
In this Letter, we use COMPAS to explore the formation of GW230529, which is a mgBHNS merger between a $3.6^{+0.7}_{-1.2}M_{\odot}$ BH and a $1.43^{+0.59}_{-0.19}M_{\odot}$ NS reported by the LVK Collaboration. By adopting the ‘delayed’ SN prescription, our population synthesis simulations can simultaneously match the inferred event rate densities of GW230529-like mgBHNS and the total BHNS mergers, along with the population distribution of the BH mass in BHNS mergers modeled by the LVK Collaboration. The mass posterior of GW230529 significantly overlaps with the highest probability region of the simulated BHNS population. Thus, we suggest that GW230529 could originate from the isolated binary evolution channel.
@@
By considering two specific EoSs of AP4 and DD2, the probabilities that GW230529 can have tidal disruption are $12.8\%$ and $63.2\%$, respectively. If GW230529 is a disrupted event, the associated kilonova is predicted to have an apparent magnitude of $\sim 23-24\,{\rm{mag}}$, and hence, can be detected by the present survey projects and the LSST in the future. Since GW230529 could be an off-axis event inferred from the GW observation, its associated GRB might be too dim to be observed by $\gamma$-ray detectors, interpreting the lack of a detected GRB associated with the event.
@@
The discovery of GW230529 revealed that the mass gap between $\sim 2.2-5\,M_{\odot}$ may not exist. Previous studies indicated that most cosmological BHNS mergers involving BH masses of $\gtrsim 5\,M_{\odot}$ are usually hard to make tidal disruption and to generate bright EM signals (e.g., Zhu et al., 2021a, b, 2022b; Fragione, 2021; Drozda et al., 2022). The existence of mgBHNS mergers suggests that BHNS mergers are still likely to be multimessenger sources that emit GWs, GRBs, and kilonovae. Although mgBHNS mergers account for $\sim 60\%$ of the cosmological BHNS population, we find that $100\%$ and $\gtrsim 90\%$ disrupted BHNS mergers are expected to originate from mgBHNS mergers, when considering the EoSs of AP4 and DD2, respectively.
@@
We thank Paul D. Lasky, Liang-Duan Liu, Ilya Mandel, Eric Thrane, Xiangyu Ivy Wang, and Yun-Wei Yu for helpful discussions. JPZ thanks the COMPAS group at Monash University. YK and LS were supported by the Beijing Natural Science Foundation (1242018), the National SKA Program of China (2020SKA0120300), and the Max Planck Partner Group Program funded by the Max Planck Society. YQ acknowledges the support of the Key Laboratory for Relativistic Astrophysics at Guangxi University.

### v2

@@ 3.1 Tidal Disruption Probability
Figure 3 shows the parameter space where the NS can be tidally disrupted using Equation (1). For a given $\chi_{{\rm BH},z}$, BHNS mergers composed of a low-mass BH component and a low-mass NS component are more easily to have tidal disruption. The mass space that allows NS tidal disruption expands significantly with increasing $\chi_{{\rm BH},z}$ and the adoption of a stiffer EoS. We display the 90% credible posterior distributions and the medians of component masses, as well as the medians of BH aligned-spins $\chi_{{\rm BH},z}$, for GW200105, GW200115, and GW230529 (Abbott et al., 2021a; The LIGO Scientific Collaboration et al., 2024) in Figure 3. Based on the GW observed samples reported by the LVK Collaboration (Abbott et al., 2021a; The LIGO Scientific Collaboration et al., 2024), the explicit results of the tidal disruption probabilities for these three BHNS mergers are listed in Table 2. It is obvious that GW200105 is unlikely to make tidal disruption to generate any EM counterparts, because its BH and NS components are massive with the mass distribution significantly outside the tidal disruption region. Although GW200115 has BH and NS masses much lighter than those of GW200105, tidal disruption can occur with only a very low probability of $P_{\rm tidal}=2.76\%$ under the EoS of DD2. Compared with GW200105 and GW200115 whose BH masses are $\gtrsim 5\,M_{\odot}$, GW230529, as a mgBHNS merger, occupies a larger portion of its mass parameter space located within the disruption region, especially when considering a stiffer EoS. More specifically, the tidal disruption probabilities are $12.8\%$ and $63.2\%$ by adopting the EoS of AP4 and DD2, respectively. The former value is similar to that reported in The LIGO Scientific Collaboration et al. (2024), which was obtained by marginalizing over the EoS.
@@ 3.4 Implication of EM detectability for BHNS Mergers
Total BHNS $17.1\%$ $0.71$ $0.01$ $0.15$ $0.70$ $0.12$ $0.18$
DD2 Mass-gap BHNS $24.9\%$ $1.04$ $0.06$ $0.79$ $1.03$ $0.20$ $0.33$
High-mass BHNS $3.4\%$ $0.14$ $0.01$ $0.05$ $0.14$ $0.02$ $0.02$
Total BHNS $28.3\%$ $1.18$ $0.27$ $0.07$ $1.17$ $0.20$ $0.35$
@@
Based on our population synthesis simulation results, we then simulate the tidal disruption probability, kilonova detectable rate, and GRB detection rate for GW BHNS mergers detected in the near future GW observing era (e.g., O4b and O5). We conservatively consider BHNS mergers occurring within $300\,{\rm Mpc}$, which is the observed distance of GW200105 and GW200115, as well as the representative distance of detectable BHNS mergers in O4 (Abbott et al., 2018b; Zhu et al., 2021b; Colombo et al., 2023; Gupta et al., 2023). We find that $100\%$ and $\gtrsim 90\%$ disrupted BHNS mergers are expected to originate from mgBHNS mergers by adopting the EoS of AP4 and DD2, respectively. When the detection depth exceeds $24\,{\rm mag}$, the survey projects can cover the majority of kilonovae associated with BHNS mergers. However, considering the rapid rise and decline of kilonovae, their signals may be challenging to be identified. Additionally, we find that the detection rate of GRBs originating from mgBHNS mergers is $\sim 0.1-0.4\,{\rm yr}^{-1}$. Therefore, in the future, it may be possible to search for kilonova signals via the target-of-opportunity follow-up observations of both GW and GRB triggers, like the multimessenger observations between GW170817/GRB170817A/AT2017gfo.
@@ 4 Conclusion
In this Letter, we use COMPAS to explore the formation of GW230529, which is a mgBHNS merger between a $3.6^{+0.7}_{-1.2}M_{\odot}$ BH and a $1.43^{+0.59}_{-0.19}M_{\odot}$ NS reported by the LVK Collaboration. By adopting the ‘delayed’ SN prescription, our population synthesis simulations can simultaneously match the inferred event rate densities of the mgBHNS and total BHNS mergers obtained from the population analyses, along with the population distribution of the BH mass in BHNS mergers modeled by the LVK Collaboration. The mass posterior of GW230529 significantly overlaps with the highest probability region of the simulated BHNS population. Furthermore, mgBHNS mergers can also originate from dynamical formation in dense stellar environments or triple system, but the expected rate densities could be much lower than inferred GW mgBHNS rate densities (e.g., Fragione et al., 2020; Gupta et al., 2020; Ye et al., 2020; Tagawa et al., 2021). Since GW230529 contributes significantly to the current inferred mgBHNS rate densities, we suggest that GW230529 could originate from the isolated binary evolution channel.
@@
By considering two specific EoSs of AP4 and DD2, the probabilities that GW230529 can have tidal disruption are $12.8\%$ and $63.2\%$, respectively. If GW230529 is a disrupted event, the associated kilonova is predicted to have an apparent magnitude of $\sim 23-24\,{\rm{mag}}$, and hence, can be detected by the present survey projects and the LSST in the future. Since GW230529 could be an off-axis event inferred from the GW observation, its associated GRB might be too dim to be observed by $\gamma$-ray detectors, interpreting the lack of a detected GRB associated with the event. Our results are generally consistent with those suggested by other recent studies (e.g., Chandra et al., 2024; Ronchini et al., 2024).
@@
The discovery of GW230529 revealed that the mass gap between $\sim 2.2-5\,M_{\odot}$ may not exist. Previous studies indicated that most cosmological BHNS mergers involving BH masses of $\gtrsim 5\,M_{\odot}$ are usually hard to make tidal disruption and to generate bright EM signals (e.g., Zhu et al., 2021a, b, 2022b; Fragione, 2021; Drozda et al., 2022). The existence of mgBHNS mergers suggests that BHNS mergers are still likely to be multimessenger sources that emit GWs, GRBs, and kilonovae. Although mgBHNS mergers account for $\sim 50\%$ of the cosmological BHNS population, we find that $100\%$ and $\gtrsim 90\%$ disrupted BHNS mergers are expected to originate from mgBHNS mergers, when considering the EoSs of AP4 and DD2, respectively.
@@
Population Synthesis Simulations in this Letter made use of the COMPAS rapid binary population synthesis code (version 02.39.00), which is freely available at http://github.com/TeamCOMPAS/COMPAS. We thank the GWOSC (Gravitational Wave Open Science Center; https://gwosc.org) for making the data of GW230529 publicly available at https://doi.org/10.7935/6k89-7q62. We thank Floor Broekgaarden for publicly providing access to a template of a table that can summarize the binary population synthesis settings (https://github.com/FloorBroekgaarden/templateForTableBPSsettings). We thank Floor Broekgaarden, Paul D. Lasky, Liang-Duan Liu, Ilya Mandel, Eric Thrane, Xiangyu Ivy Wang, Zepei Xing, and Yun-Wei Yu for helpful discussions. JPZ thanks the COMPAS group at Monash University. YK and LS were supported by the Beijing Natural Science Foundation (1242018), the National SKA Program of China (2020SKA0120300), and the Max Planck Partner Group Program funded by the Max Planck Society. YQ acknowledges the support of Anhui Provincial Natural Science Foundation (grant No. 2308085MA29). We summarize the binary population synthesis settings of our COMPAS simulations in Table 4.
Table 4: Initial values and default settings of the population synthesis simulation with COMPAS
Description and name Value/range Note/setting
Initial conditions
Initial primary mass $M_{1,\rm i}$ $[5,150]\,M_{\odot}$ Kroupa (2001) IMF $\propto{M_{1,\rm i}}^{-\alpha_{\rm IMF}}$ with $\alpha_{\rm{IMF}}=2.3$ for stars above $5\,M_{\odot}$
Initial mass ratio $q_{\rm i}=M_{2,\rm i}/M_{1,\rm i}$ $[0,1]$ We assume a flat mass ratio distribution $p(q_{\rm i})\propto 1$ with $M_{2,\rm i}\geq 0.1\,M_{\odot}$
Initial semi-major axis $a_{\rm i}$ $[0.01,1000]\,{\rm AU}$ Distributed flat-in-$\log p(a_{\rm i})\propto 1/{a_{\rm i}}$
Initial metallicity $Z_{\rm i}$ $[0.0001,0.03]$ Distributed using a uniform grid in $\log(Z_{\rm i})$ with 10 metallicities
Initial orbital eccentricity $e_{\rm i}$ 0 All binaries are assumed to be circular at birth
Fiducial parameter settings:
Stellar winds for hydrogen rich stars Belczynski et al. (2010a) Based on Vink & de Koter (2005), including LBV wind mass loss with $f_{\rm{LBV}}=1.5$
Stellar winds for helium stars Belczynski et al. (2010b) Based on Hamann & Koesterke (1998) and Vink & de Koter (2005)
Max transfer stability criteria $\zeta$-prescription Based on Vigna-Gómez et al. (2018) and references therein
Mass transfer accretion rate thermal timescale Limited by thermal timescale for stars Vink & de Koter (2005); Vinciguerra et al. (2020)
Eddington-limited Accretion rate is Eddington-limit for compact objects
Non-conservative mass loss isotropic re-emission Massevitch & Yungelson (1975); Bhattacharya & van den Heuvel (1991)
Soberman et al. (1997); Tauris & van den Heuvel (2006)
Case BB mass transfer stability always stable Based on Tauris et al. (2015, 2017); Vigna-Gómez et al. (2018)
CE prescription $\alpha_{\rm CE}-\lambda$ Based on Webbink (1984); de Kool (1990)
$\alpha_{\rm CE}$-parameter 1, 2, 5, 10
CE $\lambda$-parameter $\lambda$ Based on Xu & Li (2010) and Dominik et al. (2012)
Hertzsprung gap (HG) donor in CE pessimistic Defined in Dominik et al. (2012): HG donors don’t survive a CE phase
SN natal kick magnitude $v_{\rm k}$ $[0,\infty)\,{\rm km}\,{\rm s}^{-1}$ Drawn from Maxwellian distribution with standard deviation $\sigma_{\rm{rms}}$
SN natal kick polar angle $\theta_{\rm k}$ $[0,\pi]$ $p(\theta_{\rm k})=\sin(\theta_{\rm k})/2$
SN natal kick azimuthal angle $\phi_{\rm k}$ $[0,2\pi]$ Uniform $p(\phi)=1/2\pi$
SN mean anomaly of the orbit $[0,2\pi]$ Uniformly distributed
CCSN remnant mass prescription delayed From Fryer et al. (2012), which has no lower BH mass gap
USSN remnant mass prescription delayed From Fryer et al. (2012)
ECSN remnant mass presciption $m_{\rm{f}}=1.26\,M_{\odot}$ Based on Equation (8) in Timmes et al. (1996)
CCSN velocity dispersion $\sigma_{\rm{rms}}$ 100, $100,\,265\,{\rm km}\,{\rm s}^{-1}$ 1D rms value based on Hobbs et al. (2005) and Atri et al. (2019)
USSN and ECSN velocity dispersion $\sigma_{\rm{rms}}$ $30\,{\rm km}\,{\rm s}^{-1}$ 1D rms value based on e.g., Pfahl et al. (2002); Podsiadlowski et al. (2004)
PISN/PPISN remnant mass prescription Marchant et al. (2019) As implemented in Marchant et al. (2019)
Maximum NS mass $M_{\rm NS,max}=2.2\,M_{\odot}$ Based on Antoniadis et al. (2013); Alsing et al. (2018); Romani et al. (2022)
Tides and rotation We do not include prescriptions for tides and/or rotation
Simulation settings
Total number of binaries sampled per metallicity $10^{6}$ A million binaries per $Z_{\rm{i}}$ grid point are simulated
Binary fraction $f_{\rm{bin}}=1$
Solar metallicity $Z_{\odot}$ $Z_{\odot}$ = 0.0142 Based on Asplund et al. (2009)
Binary population synthesis code COMPAS (v02.39.00) Stevenson et al. (2017); Vigna-Gómez et al. (2018); Neijssel et al. (2019)
Broekgaarden et al. (2019); Team COMPAS: Riley, J. et al. (2022)